OpenStudy (anonymous):

4/3 + 2/3x = 13/6 + 5/2x + 1/2

OpenStudy (anonymous):

to simplify your equation you should multiply all numbers with 6, then you've got: 8+4x=13+15x+3 then you have to solve the term with arithmetics ^^ simplify the term again: 8+4x=16+15x minus 4x and 16 -8=11x x= -8/11 well maybe i got something wrong ^^